# Beyond Subspace Isolation: Many-to-Many Transformer for Light Field Image Super-resolution

### v1

@@ IV-C Qualitative Comparisons
We further explore the superior performance of LF-SANet in qualitative evaluation. Figure 4 presents qualitative results at the 4×4$\times$4 × scale for three representative samples, namely (a) Perforated_Metal_3, (b) Palais_du_Luxembourg and (c) Bicycle. The first two samples are from the EPFL dataset captured by Lytro cameras [19], and the third one is from the synthetic dataset HCInew. We compared M2MT-Net with four other methods: IINet and DistgSSR represent the convolution-based methods, while LFT and EPIT are from the Transformer-based group. Zoom-in views inside blue and red boxes are provided to show more details. Accompanying these visuals, PSNR/SSIM calculated on the red box areas. In general, all these techniques capably enhance resolution and preserve primary structures, but nuanced distinctions emerge within the details, especially the zoom-in views of red boxes.
In the Perforated_Metal_3 sample, most methods portray the perforated hole reasonably well but fall short in edge sharpness, likely influenced by lighting and occlusion challenges. M2MT-Net, however, produces notably sharper edges and more round shape of the hole. In the Palais_du_Luxembourg sample, M2MT-Net excels in reconstructing the textures of the bricks beyond the other methods. For the Bicycle sample, the edges of the books are clearly sharper in M2MT-Net than others.
In Figure 5, SAI-wise PSNR on these three samples is visualized. The visual representation highlights M2MT-Net’s notable enhancements across SAIs. Notably, in cases (a) Perforated_Metal_3 and (c) Bicycle, the lowest PSNR values achieved by M2MT-Net are still higher than the highest PSNR values of other methods. This signifies the consistent superiority of M2MT-Net.
@@ IV-D LAM Analysis
To further probe into the underlying capability of M2MT-Net, we employ the Local Attribution Map (LAM) technique [33] to discern influential pixels, thereby providing insight and transparency into the performance of M2MT-Net. In Figure 4, the LAM visualization is provided below PSNR/SSIM. It is calculated on the blue box regions with the red box regions as targets. Diffusion Index (DI) accompanies the visualization as a qualitative indicator of pixel utilization.
From the results, M2MT-Net’s superior utilization across the three samples is evident, demonstrating significantly more active pixels both within and across SAIs. This superiority is further supported by the DI values as M2MT-Net consistently records values above 20, while competing methods hover at 10 or even lower. Delving deeper into these active pixels reveals intriguing insights.
@@
For instance, in the Perforated_Metal_3 sample, though repeated perforated holes offer potential patterns for reconstruction, most methods focus solely on the neighboring area. In contrast, M2MT-Net’s active pixels span the entire column, suggesting that it identifies shared characteristics among the holes in the column and leverages them as complementary cues. Similarly, in the Palais_du_Luxembourg sample, the texture of brick walls exhibit recurring patterns for reuse. M2MT-Net manages to utilize not only the specific brick wall but also the neighboring one on the right, and the influential pixels have high activities across SAIs, hence the patterns are recovered with visible edges, unlike its counterparts generate a blurry area due to their narrower focus and weaker correlation across SAIs. For the Bicycle sample, the books on the shelves present similar patterns as well. M2MT-Net’s advantage becomes evident as it activates pixels not only around the local area but also on a book in a similar red color on the shelf above, regardless of the distance.
The DI values offer insights into the relationship between model performance and pixel utilization. In theory, higher DI indicates better pixel utilization and should result in improved performance. This holds true for M2MT-Net, as it achieves a double-digit DI alongside significantly higher PSNR and SSIM scores, and it still holds true if comparing only within the convolution-based or Transformer-based groups.
@@
The DI values shed light on the relation between model performance and pixel utilization. In theory, higher DI indicates higher pixel utilization and should result in better performance. It holds true for M2MT-Net as having a double-digit DI and a significantly higher PSNR and SSIM, and it remains consistent when comparing only within the convolution-based or Transformer-based groups. However, when comparing across these two groups, a different trend emerges as convolution-based IINet and DistgSSR tend to have higher DI values but often lag behind Transformer-based LFT and EPIT in terms of PSNR and SSIM. This highlights the distinct nature of pixel utilization between convolutions and Transformers, where convolutions leverage more pixels but are constrained by locality, while Transformers use fewer pixels in a more effective way through its global dependency modeling mechanism.
@@
Perforated_Metal_3 Ground-truth IINet [50] DistgSSR [22] LFT [31] EPIT [32] M2MT-Net
MSE×$\times 100$ 0.968 0.832 1.037 0.918 0.706
@@
Sphynx Ground-truth IINet [50] DistgSSR [22] LFT [31] EPIT [32] M2MT-Net
MSE×$\times 100$ 0.178 0.139 0.149 0.147 0.136
@@
bicycle Ground-truth IINet [50] DistgSSR [22] LFT [31] EPIT [32] M2MT-Net
MSE×$\times 100$ 2.805 2.633 2.652 2.557 2.324
@@
monasRoom Ground-truth IINet [50] DistgSSR [22] LFT [31] EPIT [32] M2MT-Net
MSE×$\times 100$ 0.584 0.571 0.590 0.567 0.560
@@ IV-E Angular Consistency
While the reconstruction of visual detail is important for LFSR, the preservation of parallax structure within LF images is equally crucial. This aspect of parallax structure cannot be adequately discerned solely by examining the reconstructed SAIs. Thus, to comprehensively assess the angular consistency, we conduct an evaluation through depth estimation. OACC-Net [66] is applied to generate depth maps on the super-resolved output of the methods under comparison. The depth estimated from HR images serves as the ground truth for this evaluation. Figure 6 visually represents the depth maps for two real-world and synthetic examples, accompanied by the $MSE\times 100$ as a quantitative evaluation metric.
M2MT-Net’s superiority, as highlighted in Perforated_Metal_3 of Figure 4, is corroborated in the generated depth map. This method successfully reconstructs perforated holes with crisply integrated edges spanning from near to distant from the camera as evidenced in the blue and red boxes. In stark contrast, competing methods struggle, yielding blurred and entangled edges in this complex scene. When examining scenes featuring salient objects, M2MT-Net continues to excel. In the Sphynx sample, the contours of the sphynx’s nose, mouth and neck are distinctly delineated in M2MT-Net’s depth map. Other methods, however, generate noticeable blurs or artifacts in these areas. The bicycle sample further illustrates M2MT-Net’s proficiency, where distinct separations between the bicycle’s handlebar and the background are evident, as well as more continuous structures in the plant’s trunks. Other methods falter, blending the handlebar’s contour with the background or breaking the trunk’s structure into fragments. Finally, in the monasRoom example, M2MT-Net’s depth map reveals a smoother surface on the T-shaped object and sharper contours on the leaf with fewer inaccuracies, demonstrating a closer approximation to the ground-truth when compared to the other methods, which produce noticeable bumpy artifacts.
@@
TABLE II: Comparison of computational efficiency against the state-of-the-art methods at the 4×4$\times$4 × scale. Runtime is the average elapsed time per sample over all 23 samples. #Params is the number of trainable parameters.
Method Type Runtime/s #Params
LF-IINet [50] Convolution-based 1.34 4,885,824
DistgSSR [22] Convolution-based 1.70 3,581,568
LFT [31] Transformer-based 5.84 1,163,392
EPIT [32] Transformer-based 3.16 1,470,080
M2MT (Ours) Transformer-based 2.03 3,986,400
(a) PSNR vs Runtime
(b) PSNR vs Parameter Number
@@
Figure 7: Tradeoff between performance and efficiency at the 4×4$\times$4 × scale. The convolution-based methods are in red while the Transformer-based methods are in green. Candidates positioned closer to the top-right corner of the plots are better.
IV-F Computational Efficiency
We evaluate the computational efficiency of M2MT-Net against the best competitors LF-IINet, DistgSSR, LFT, and EPIT. Two metrics are involved: the running speed and the number of trainable parameters. The running speed measures the elapsed time per sample when running all 23 samples. The result is shown in Table II, and the tradeoff between PSNR and runtime/parameter numbers is visualized as plots in Figure 7. The convolution-based methods are in red while the Transformer-based methods are in green. In general, candidates positioned closer to the top-right corner of the plots indicate a more favorable tradeoff between performance and efficiency.
In terms of runtime, LF-IINet and DistgSSR are the fastest method as its convolution operations are local, making it quicker than global Transformer operations. Yet, within Transformer-based techniques, M2MT-Net emerges as the swiftest, taking approximately 20% more time than DistgSSR but being 35% and 65% faster than EPIT and LFT respectively. This advantage is attributed to M2MT reducing the channel dimension from $UVC$ to $C_{Cor}$ via its linear layer in correlation encoding, significantly trimming the computational demand of the following self-attention component. On the flip side, this channel conversion makes M2MT bulkier in model size, as the linear layers require more parameters in both correlation encoding and decoding. Still, LF-SANet has 19% fewer parameters than LF-IINet and 11% more parameters than DistgSSR, keeping its size fairly manageable.

### v2

@@ IV-C Qualitative Comparisons
We further explore the superior performance of M2MT-Net in qualitative evaluation. Fig. 4 presents qualitative results at the 4×4$\times$4 × scale for three representative samples, namely (a) Perforated_Metal_3, (b) Palais_du_Luxembourg and (c) Bicycle. The first two samples are from the EPFL dataset captured by Lytro cameras [19], and the third one is from the synthetic dataset HCInew. We compared M2MT-Net with five methods: DistgSSR and HLFSR represent the convolution-based methods, while LFT, EPIT and LF-DET represent the Transformer-based methods. Zoom-in views inside blue and red boxes are provided to show more details. Accompanying these visuals, PSNR and SSIM are calculated on the red box areas. In general, all these techniques capably enhance resolution and preserve primary structures, but nuanced distinctions emerge within the details, especially the zoom-in views of red boxes.
In the Perforated_Metal_3 sample, most methods portray the perforated hole reasonably well but fall short in edge sharpness, likely influenced by lighting and occlusion challenges. M2MT-Net, however, produces notably sharper edges and a more round shape of the hole. In the Palais_du_Luxembourg sample, M2MT-Net excels in reconstructing the edges of windows beyond the other methods. For the Bicycle sample, the edges of the leaves are clearly sharper in M2MT-Net than others.
In Fig. 5, SAI-wise PSNR on these three samples is visualized. The visual representation highlights M2MT-Net’s notable enhancements across SAIs. Notably, in cases (a) Perforated_Metal_3 and (c) Bicycle, the lowest PSNR values achieved by M2MT-Net are still higher than the highest PSNR values of other methods. This observation signifies M2MT-Net’s consistent superiority across SAIs.
@@ IV-D LAM Analysis
To further probe into the underlying capability of M2MT-Net, we employ the Local Attribution Map (LAM) technique [42], an attribution approach to identify pixels in the input that have a significant impact on the generation of a target window in the output, to provide insight and transparency into the performance of M2MT-Net.
Assuming $\mathcal{F}(\cdot)$ is the super-resolution network as stated before and $\mathcal{D}(\cdot)$ is a detector of edges and textures, with the detector operating on the super-resolved result as $\mathcal{D}(\mathcal{F}(\cdot))$, the LAM is derived by calculating its path integrated gradient along a gradually changing path function $\gamma(\cdot)$ as follows:
[…]
where $i$ is the dimension index, and $m$ and $k$ are the number of steps and the step index in the path, respectively. $k$ is set to 50 in the analysis. Here, the detector $\mathcal{D}$ is a simple gradient detector of a local window located at a specified location $(x,y)$ of size $l\times l$ as:
$\mathcal{D}_{xy}(I_{LR})=\sum\nolimits_{i\in[x,x+l],j\in[y,y+l]}{\nabla_{ij}I_{LR}}.$ (14)
The path function $\gamma(\cdot)$ utilizes a Gaussian blur kernel to compute the blurred version of the input image, reducing the high-frequency components in the image to represent absent features:
$\gamma(\frac{k}{m})=\omega(\sigma-\frac{k}{m}\sigma)\otimes I_{LR}$ (15)
where $\omega(\sigma)$ is the Gaussian kernel with the kernel width $\sigma$, and $\otimes$ denotes the convolution operation. Under this definition, $\gamma(\cdot)$ returns the original LR image $I_{LR}$ when $k=m$ and the completely blurred LR image $I^{\prime}_{LR}$ when $k=0$, i.e., $\gamma(0)=I^{\prime}_{LR}$ and $\gamma(1)=I_{LR}$.
The Diffusion Index (DI) can be derived from LAM as a quantitative indicator of pixel utilization. It is calculated based on the Gini coefficient $G$ measuring the inequality of pixels’ impact:
$DI=(1-G)\times 100$ (16)
$G=\frac{\sum_{i}\sum_{j}|g_{i}-g_{j}|}{2n^{2}\bar{g}}$ (17)
where $g_{i}$ represents the LAM value of pixel $i$, $n$ is the total number of pixels, and $\bar{g}$ is the average LAM value. Essentially, a high DI value indicates a model’s capacity to involve a broader range of pixels in the generation of the target window, while a low DI value suggests a more limited involvement.
Although the LAM technique was initially developed for single image super-resolution, it can be seamlessly adapted for light field super-resolution without significant modifications because the BasicLFSR framework [70] processes a 4D LF image as a 2D macro-pixel (MacPI) image [22].
The LAM visualization and the DI are provided below PSNR/SSIM in Fig. 4. The DI is calculated on the blue box regions with the red box regions as targets. The LAM results show that M2MT-Net consistently exhibits more activated pixels both within and across SAIs. This superiority is substantiated by the DI values as M2MT-Net is the highest, ranging between 19.7059 and 25.2688. This is 5-6% higher than the second-ranked method, LF-DET, whose DI values range from 18.7618 to 23.9182. Meanwhile, the DI values of other competing methods are significantly lower, hovering from 5.1459 to 13.1903.
@@
Delving deeper into these activated pixels reveals intriguing insights. For instance, in the Perforated_Metal_3 sample, though repeated perforated holes offer potential patterns for reconstruction, most methods focus solely on the neighboring area. LF-DET has some activated pixels on distance holes; however, the activation is weak. In contrast, M2MT-Net’s activated pixels span not only the same column but also the neighboring columns with high activation, indicating that it identifies shared characteristics among the holes and leverages them as complementary cues. Similarly, in the Palais_du_Luxembourg sample, the building’s windows exhibit recurring patterns for reuse. M2MT-Net manages to utilize not only the windows in the red box but also the ones in a broader area of the blue box, and the influential pixels have high activities across SAIs. Hence, the patterns are recovered with visible edges, unlike its counterparts, which generate a blurry area due to their narrower focus and weaker correlation across SAIs. For the Bicycle sample, the plant leaves present similar patterns. M2MT-Net’s advantage becomes evident as it activates pixels on leaves not only on the same trunk but also on the other trunk.
@@
The DI values shed light on the relation between model performance and pixel utilization. In general, higher DI indicates higher pixel utilization and should result in better performance. It holds true for LF-DET and M2MT-Net as their DI values are significantly high as well as their PSNR and SSIM, and it remains consistent when comparing only within the convolution-based or Transformer-based groups. However, when comparing these two groups, a different trend emerges as a high DI does not necessarily mean high PSNR and SSIM, such as DistgSSR and LFT. This highlights the distinct nature of pixel utilization between convolutions and Transformers, where convolutions leverage more pixels but are constrained by locality, while Transformers establish long-range dependencies among broader pixels, though these dependencies may not always be strong enough to aid in super-resolution as effectively as M2MT-Net.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Perforated_Metal_3 Ground-truth DistgSSR HLFSR LFT EPIT LF-DET M2MT-Net
MSE×$\times 100$ 0.831 0.796 1.046 0.919 0.834 0.613
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Sphynx Ground-truth DistgSSR HLFSR LFT EPIT LF-DET M2MT-Net
MSE×$\times 100$ 0.136 0.150 0.149 0.146 0.165 0.111
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
bicycle Ground-truth DistgSSR HLFSR LFT EPIT LF-DET M2MT-Net
MSE×$\times 100$ 2.627 2.648 2.653 2.556 2.609 2.224
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
monasRoom Ground-truth DistgSSR HLFSR LFT EPIT LF-DET M2MT-Net
MSE×$\times 100$ 0.571 0.556 0.590 0.567 0.598 0.502
@@ IV-E Angular Consistency
While the reconstruction of visual detail is important for LFSR, the preservation of parallax structure within LF images is equally crucial. This aspect cannot be adequately discerned solely by examining the reconstructed SAIs. Thus, to comprehensively assess the angular consistency, we conduct an evaluation through depth estimation. OACC-Net [77] is applied to generate depth maps on the super-resolved output of the methods under comparison. The depth estimated from HR images serves as the ground-truth for this evaluation. Fig. 6 visually represents the depth maps for two real-world and synthetic examples, accompanied by the $MSE\times 100$ as a quantitative evaluation metric.
M2MT-Net’s superiority, as highlighted in Perforated_Metal_3 of Fig. 4, is corroborated in the generated depth map. This method successfully reconstructs more perforated holes with integrated edges spanning from near to distant from the camera as evidenced in the blue and red boxes. In stark contrast, competing methods struggle, yielding blurred and entangled edges in this complex scene. When examining scenes featuring salient objects, M2MT-Net continues to excel. In the Sphynx sample, the contours of the sphynx’s nose, mouth and neck are distinctly delineated in M2MT-Net’s depth map. Other methods, however, generate noticeable blurs or artifacts in these areas. The bicycle sample further illustrates M2MT-Net’s proficiency, where distinct separations between the bicycle’s handlebar and the background are evident, as well as more continuous structures of the plant’s trunks. Other methods falter, blending the handlebar’s contour with the background or breaking the trunk’s structure into fragments. Finally, in the monasRoom example, M2MT-Net’s depth map reveals a smoother surface on the T-shaped object and integrated shapes of the leaf with fewer holes, demonstrating a closer approximation to the ground-truth when compared to the other methods, which produce noticeable bumpy artifacts.
@@
TABLE II: Comparison of model efficiency and performance with the state-of-the-art methods by varying the number of blocks at the 4×4$\times$4 × scale. Time is the inference time. #Params. is the number of parameters. Memory is the peak GPU memory usage for training. FLOPs is the number of floating-point operations. Time is the inference time. #blocks is the number of blocks.
Method #Params. (M) Memory (GB) FLOPs (G) Time (s) PSNR/SSIM
LF-IINet [58] 4.886 1.99 57.36 1.55 29.04/0.9188
LFSSR-SAV [27] 1.542 8.25 99.45 3.20 29.37/0.9223
DistgSSR [22] 3.582 4.43 65.26 1.89 28.99/0.9195
HLFSR [62] 13.865 2.43 45.73 6.83 29.20/0.9222
LFT [40]
#blocks = 4 1.163 6.41 30.20 6.22 29.33/0.9196
#blocks = 8 2.150 11.11 55.64 12.27 29.44/0.9219
#blocks = 12 3.136 16.80 81.07 16.80 29.59/0.9238
EPIT [41]
#blocks = 4 1.212 7.25 57.87 2.25 29.20/0.9170
#blocks = 5 1.470 8.63 74.15 2.61 29.31/0.9196
#blocks = 8 2.246 12.77 110.98 3.77 29.50/0.9212
#blocks = 12 3.328 18.30 164.09 5.28 29.58/0.9212
#blocks = 14 3.797 21.06 190.65 6.12 29.53/0.9216
LF-DET [23]
#blocks = 3 1.293 13.80 39.17 3.79 29.21/0.9199
#blocks = 4 1.697 17.83 51.20 4.81 29.42/0.9220
#blocks = 5 2.080 21.86 63.23 5.91 29.47/0.9228
M2MT-Net (Ours)
#blocks = 3 1.557 2.72 14.40 1.14 29.30/0.9222
#blocks = 4 2.043 3.20 18.29 1.33 29.50/0.9226
#blocks = 5 2.529 3.67 22.18 1.52 29.51/0.9239
#blocks = 6 3.015 4.15 26.07 1.71 29.58/0.9253
#blocks = 7 3.501 4.62 29.96 1.90 29.67/0.9265
#blocks = 8 3.986 5.10 33.85 2.09 29.85/0.9284
#blocks = 9 4.472 5.57 37.74 2.28 29.74/0.9259
(a) PSNR vs Parameter Number (b) PSNR vs Peak GPU Memory Usage (c) PSNR vs FLOPs (d) PSNR vs Inference Time
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 7: Tradeoff between performance and efficiency at the 4×4$\times$4 × scale. Candidates positioned closer to the top-left corner of the plots have a better performance-efficiency tradeoff.
IV-F Model Efficiency
We evaluate the model efficiency of M2MT-Net against the top competitors, the convolution-based LF-IINet, LFSSR-SAV, DistgSSR and HLFSR, and the Transformer-based methods, LFT, EPIT and LF-DET at the 4×4$\times$4 × scale using four metrics: number of parameters, peak GPU memory usage, the floating-point operations (FLOPs), and inference time. The number of parameters and peak GPU memory reflect memory complexity, indicating the theoretical model size and the actual minimum memory required for training, respectively. FLOPs and inference time reflect computational complexity, representing the theoretical number of operations for processing a $32\times 32$ LF patch and the actual average time for inferring a sample from the test datasets with GPU acceleration. FLOPs is obtained by utilizing the fvcore library [78].
To ensure a fair and comprehensive evaluation, we train and evaluate variants of the Transformer-based methods, LFT, EPIT, LF-DET, and our M2MT-Net, varying the number of blocks (#blocks, equivalent to $n_{2}$ in M2MT-Net). The results are compiled with PSNR on the EPFL dataset in Table II. To better understand how these methods balance performance and efficiency, we also plot the four efficiency metrics on the x-axis against PSNR on the y-axis in Fig. 7. In these plots, models closer to the top-left corner represent a more favorable balance between performance and efficiency.
Regarding memory complexity, M2MT-Net is very similar to its Transformer-based peers in terms of the parameter number with minor PSNR differences under 0.1 dB among variants with similar parameter numbers in Fig. 7 (a). However, a stark contrast emerges in peak GPU memory usage as M2MT-Net’s performance-efficiency curve (represented in pink) consistently trends toward the top-left direction relative to its competitors, LFT (in green), EPIT (in blue), and LF-DET (in yellow), in Fig. 7 (b). Specifically, M2MT-Net variants require less than 5.6 GB of GPU memory, whereas LFT starts at 6.41 GB, EPIT at 7.25 GB, and LF-DET at a hefty 13.80 GB. Notably, the 5-block LF-DET variant, while achieving a similar PSNR to the 4-block M2MT-Net (29.47 dB vs. 29.50 dB), consumes a significant 21.86 GB GPU memory. This is over six times the GPU memory used by the 4-block M2MT-Net (3.20 GB) and nearly maxes out the 24 GB memory of an Nvidia GTX 3090 GPU. LF-DET’s high memory demand is primarily due to its complex design, which incorporates two spatial Transformers and three angular Transformers per block, which results in a memory bottleneck, significantly limiting LF-DET’s scalability compared to its competitors. In contrast, M2MT-Net maintains low memory usage due to its streamlined and compact design.
From a computational complexity standpoint, M2MT-Net exhibits an exceptional performance-efficiency tradeoff in terms of FLOPs and inference time as its performance-efficiency curve consistently positions itself in the top-left direction relative to other methods’ curves in Fig. 7 (c) and (d). The 6-block M2MT-Net, with a PSNR of 29.58 dB, matches or surpasses the best-performing variants of the other models, such as the 12-block LFT and EPIT (29.59 dB and 29.58 dB, respectively) and the 5-block LF-DET (29.47 dB). Remarkably, it requires only up to 41.23% of the FLOPs (26.07 G) and 32.39% of the inference time (1.71s) compared to the most efficient variants of these competitors (63.23 G by the 5-block LF-DET and 5.28s by the 12-block EPIT). Our top-performing 8-block M2MT-Net exceeds other methods by more than 0.26 dB in PSNR while still demanding lower FLOPs and inference time (33.85 G and 2.09s) than the most lightweight variants of nearly all other Transformer-based methods. On the other hand, while the PSNR of LFT demonstrates an upward trend with the addition of more blocks, its scalability is severely constrained due to its excessive inference time. Specifically, the 12-block variant of LFT requires an unmanageable inference time of 16.80 seconds, more than double that of its slowest competitor, the 14-block EPIT. This prohibitive inference duration renders further scaling of LFT impractical.
Meanwhile, the convolution-based methods generally exhibit a weaker performance-efficiency tradeoff as they are positioned lower and to the right compared to the Transformer-based methods, except LFSSR-SAV has a position similar to EPIT and LFT on the graph. Due to the inherent requirements of convolutional kernels, many convolution-based methods such as LF-IINet, DistgSSR, and HLFSR necessitate a larger number of parameters. This trend is illustrated in Fig. 7 (a). Notably, HLFSR requires more than 13.865 million parameters, which exceeds the graph’s range.
In summary, these results demonstrate that M2MT-Net achieves excellent LFSR performance-efficiency balance and model scalability, making it a highly favorable choice for practical LFSR applications.